OpenStudy (anonymous):

What is the solution set of 2x^2 - 3x = 5

OpenStudy (dumbcow):

use quadratic formula

OpenStudy (dumbcow):

or try to factor, this one factors

jhonyy9 (jhonyy9):

3 +/- (7) x_1,_2 = -------- = ---- x_1 = 10/4 = 5/2 4 ---- x_2 = -4/4 = -1